If twice the smaller of two consecutive numbers is added to the larger, the result is 7. Find the smaller one.
Answer by RAY100(1637)

Let x = smaller no
Let (x+1) = larger no
2S+L=7
2x+(x+1)=7
3x+1=7
3x=6
x=2
smaller =2
larger=3